OpenStudy (anonymous):

Evaluate the indefinite integral of: (x+4)/(x^2 +8x +17) The denominator can be simplified to ((x+4)^2 +1) by completing the square. You probably need to use partial fraction decomposition.

OpenStudy (anonymous):

 u-substiution twice with u=x+4 and s=u^2+1